how many atoms of FE are needed to react with 50 grams of AgNO3


1
Expert's answer
2021-06-02T06:03:06-0400

Fe + AgNO3 -->> FeNO3 + Ag

Molar Mass of Fe = 55.845

Molar Mass of AgNO3 = 169.87

= 50/169.87 = 0.294

= 0.294 × 6.023 × 1023

= 1.771 × 1023 Atoms